a paddle wheeler goes 10 miles against the current in a river in the same time that it goes 15 miles with the current. If the speed of the current is 3 mph, find the speed of the boat in the water.
Answer by mananth(16946)


let speed of boat be x mph
speed of current = 3mph
..speed upstream
x-3 mph
distance =10
time upstream = 10/(x-3)
..
downstream
speed = x+3
distance = 15 miles
time = 15/(x+y)
..
times are same
10/(x-3)=15/(x+3)
10(x+3)=15(x-3)
10x+30=15x-45
15x-10x=45-30
5x=15
x= 3 mph.... speed of boat
...
m.ananth@hotmail.ca